Q = 12000 - 60P




C (Q) = 10000 + 4Q




What is the company's profit function?

1
Expert's answer
2022-05-02T16:10:22-0400

What is the company's profit function in terms P?

(Q) = 10 000 + 4Q = 10000 + 4(12000 - 60 P) = 58000 - 240 P

Profit = TR - Q= 12 000 P - 60 P2 - (58 000 - 240 P) = -60 P2 + 12240P - 58 000